The Ricci equation is often used to estimate pKa values of oxoacids. The equation is: pKa = 8 - 9f + 4(n-m)

for an acid with the formula HmXOn, with f = formal charge on X IF ALL OXYGEN ATOMS MAKE SINGLE BONDS TO X.

Use the Ricci equation to estimate the pKa of each of the following acids: H2SeO4, H2SeO3, HSeO3-, HNO3, H3BO3.
Answer #1

H₂ se on m22, nay f = 2 Ise- o- on Pla = 8-9(2) + 464-2) - 8-18 & 4 (2) = 8-18 + 8 - -2 H₂ se ₃ m=2 na 3 f=1 Pk. 8-9(12 +4 (3यार Hse og ó - sen mar n=3 S-1 Pla= 8-9111 +4(3-1) = 8-9 + 8 = 169 - 7 HNO3 AZ_ m=1 n=3 f=2 Pka = 8-9 (21+4 (3-1) . - 8-18 +